A postal clerk sold 50 stamps for $15.84. Some were 37 cent stamps, and the rest were 23 cent stamps. Find the number of 37 cent stamps that were sold.
Answer by stanbon(75887)

A postal clerk sold 50 stamps for $15.84. Some were 37 cent stamps, and the rest were 23 cent stamps. Find the number of 37 cent stamps that were sold.
---------------
Equations::
value + value = value
37x + 23(50-x) = 1584 cents
---------------
37x + 23*50 - 23x = 1584
------
14x = 434
-----
x = 31 (# of 37 cent stamps)
50-x = 29 (# of 23 stamps)
=========================
Cheers,
Stan H.
=================